OpenStudy (anonymous):

The expression below is not a polynomial. A). Factor out the variable with the lowest or most negative exponent from the express. B) Factor completely. x^-3-2x^-4-8x^-5. Factor out x^-5

OpenStudy (anonymous):

@sauravshakya

OpenStudy (anonymous):

(x^-5)(x^2-2x-8)

OpenStudy (anonymous):

ok so is that what im looking for in a part?

OpenStudy (anonymous):

or do i fact

OpenStudy (anonymous):

or that now?

OpenStudy (anonymous):

that is not completely factored. that is just factoring out x^-5 the completely factored equation is (x^-5)(x-4)(x+2)

OpenStudy (anonymous):

but that would be my B answer right?

OpenStudy (anonymous):

if B is factoring out completely, than B is my second expression

OpenStudy (anonymous):

the (x^-5)(x-4)(x+2)

OpenStudy (anonymous):

can u help me on another one?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

im going to post it give me just a sec.

OpenStudy (anonymous):

just write it here

OpenStudy (anonymous):

or not

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The carters can purchase Season's choice blend for 1.81 per pound and Garden Mix blend for 1.21 per pound. If they wish to purchase 20 pounds and spend 32 dollars on birdseed, how many pounds of each type should they buy.

OpenStudy (anonymous):

It has 2 parts. The 1st part ask how many pounds of season's choice blend should they buy

OpenStudy (anonymous):

start buy making a system of equations. Say x is the number of pounds of Season's choice blend and y is the number of pounds of Garden Mix. how would you set up the equation?

OpenStudy (anonymous):

x+y maybe

OpenStudy (anonymous):

not to sure

OpenStudy (anonymous):

you're getting there. x+y would be the total pounds they are buying and they plan to buy 20 pounds so x+y=20

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now whenever you have a system of equations, you need the same number of equations as there are unknown variables. In this case, you don't know x and y so you need 2 equations

OpenStudy (anonymous):

the second one would be 1.81x + 1.21y = 32 or the total cost

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

there are two ways to solve this. I'll just show you substitution. solve one equation for on e variable and input it into the other one. so x=20-y. put that into 1.81x + 1.21y = 32 and solve for y. 1.81(20-y) + 1.21y = 32

OpenStudy (anonymous):

try solving for y

OpenStudy (anonymous):

7 right?

OpenStudy (anonymous):

so would that be the answer for the Season's choice or Garden mix

OpenStudy (anonymous):

yes. y is the pounds of Garden Mix.

OpenStudy (anonymous):

put 7 into x+y=20 for y and solve for x

OpenStudy (anonymous):

what about the season's choice?

OpenStudy (anonymous):

x is season's choice

OpenStudy (anonymous):

so 13

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

may i ask u another one right here?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

find a expression in factored form, for the perimeter of the figure.